OpenStudy (anonymous):

Quadratics: For each of the following graphs, state the co-ordinates of the turning point and whether it is a maximum or a minimum. a) y=(x+3)²-7 b)y=-(x-4)²+2

OpenStudy (anonymous):

What's the vertex of a?

OpenStudy (anonymous):

I don't know...

OpenStudy (anonymous):

Vertex form is y=a(x-h)^2+k the vertex being (h,k) so you should easily be able to see what the vertex is. it's (-3,-7) right?

OpenStudy (anonymous):

I have no idea.... I just got this worksheet from my teacher who terrifies me and I don't want to talk to him.

OpenStudy (anonymous):

I don't understand how to do it at all.

OpenStudy (anonymous):

Just use him for his wisdom. y=a(x-h)^2+k looks alot like y=(x+3)²-7 right? what im saying is when it's in this form you can look at it and determine the vertex (h,k) which will give you the maximum or minimum, depending on which way the graph opens. So the vertex of this one would be (-3,-7)

OpenStudy (anonymous):

What's a vertex?

OpenStudy (anonymous):

There is no negative in front of the parentheses, which means it opens up, so there will be a minimum value